OpenStudy (briana.img):

Find the height of a rectangular prism if the surface area is 458.17in^2, the width is 6.2in and the length is 8.3in.

OpenStudy (anonymous):

Find the surface area of all the sides with a width and length first can you do that?

OpenStudy (anonymous):

? hello?

OpenStudy (briana.img):

@Icedragon50 trying to figure out if i can lol

OpenStudy (anonymous):

ok

OpenStudy (briana.img):

@Icedragon50 just got one surface area, 8.3(6.2)=51.46

OpenStudy (anonymous):

well heres a hint: two sides have an area of width * length

OpenStudy (briana.img):

@Icedragon50 yeah i did that and got 51.46 lol

OpenStudy (anonymous):

Very good! now since theres TWO sides so multiply by 2 to get:102.96

OpenStudy (anonymous):

102.92 srry

OpenStudy (briana.img):

@Icedragon50 aaah okay so what's next?

OpenStudy (anonymous):

Ok so now we have 4 more faces. 2 of them are width*height and length*height

OpenStudy (anonymous):

width*Height would be 6.2*h=6.2h

OpenStudy (anonymous):

length*height would be 8.3*h=8.3h

OpenStudy (anonymous):

now you solve for h (height) by using this equation: 102.92+6.2h+8.3h=458.17

OpenStudy (briana.img):

@Icedragon50 h=24.5?

OpenStudy (anonymous):

102.92+14.5h=458.17 14.5h=458.17-102.92 14.5h=355.25 h=24.5 So, height is 24.5

OpenStudy (anonymous):

yes lol

OpenStudy (briana.img):

@Icedragon50 that isn't any of my options :(

OpenStudy (anonymous):

wit are your options??

OpenStudy (briana.img):

@Icedragon50 10.25 12.25 14.25 16.25

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

hmmm gimmee a sec

OpenStudy (nincompoop):

do you know what surface area is?

OpenStudy (nincompoop):

first, draw a rectangular prism and see if you can better visualize the problem

OpenStudy (briana.img):

@nincompoop i did on my paper

OpenStudy (nincompoop):

draw it here

OpenStudy (anonymous):

OK THIS TOOK ME FOREVER BUT HERE IT IS

OpenStudy (anonymous):

SA = Surface Area = 458.17 in² L = Length = 8.3 in W = Width = 6.2 in H = Height = What we're solving for. The formula for the surface area of a rectangular prism is SA = 2*(LW + LH + WH) Substituting values: 458.17 = 2*(8.3*6.2 + 8.3H + 6.2H) 229.085 = 51.46 + 14.5H 14.5H = 229.085 - 51.46 H = 117.625/14.5 H = 12.25 inches. That's about it! =)

OpenStudy (briana.img):